OpenStudy (anonymous):

what is derivative of 1+ 2/X+ 3/X+4

OpenStudy (anonymous):

d/dx(1 + 2/x + 3/x + 4) = d/dx(1) + d/dx(2/x) + d/dx(3/x) + d/dx(4) = 0 + (-1)(2)(1/x^2) + (-1)(3)(1/x^2) + 0 -2/x^2 -3/x^2

OpenStudy (anonymous):

\[\huge -5^{-2}\]

OpenStudy (anonymous):

??